A 2.24L cylinder of oxygen at N.T.P is found to develop a leakage. When the leakage was plugged the pressure dropped to 570mmof Hg. The number of moles of gas that escaped will be-
A) 0.025
B) 0.050
C) 0.075
D) 0.09

Grade:12

1 Answers

Arun
5 years ago
Original pressure of Hg= 760
drop in pressure= 570
change in pressure= 760-570= 190mm
= 190/760 atm

PV= nRT
n = P V/ RT
= 190 x 2.24/ 760 x 0.821 x 273 
= 0.0249 moles